Two vectors are given: 7 = 3 + 27, y = 27 - 8j.

a) Draw these vectors in Cartesian coordinate system

b) Represent both of these vectors in magnitude, direction notation.

C) Represent vector 27 + 3ÿ in magnitude, direction notation.


1
Expert's answer
2021-11-07T19:26:06-0500

1) Draw:



2)

"\vec r_1=\sqrt{3^2+2^2}\angle\arctan\frac{2}{3}=3.61\angle33.7°,\\
\vec r_2=8.25\angle283.25°."

3)

"|r|=\sqrt{27^2+3^2}=27.2,\\\space\\
\theta=\arctan\frac{3}{27}=6.3°"